OpenStudy (anonymous):

the statement 2^n=(nC0)+(nC1)+(nC2)+...+(nCn). 1). Prove the statement is true for all n>=0 by induction on n. [(nC1)means n choose 1]

OpenStudy (anonymous):

\[2^{n}=\left(\begin{matrix}n \\ 0\end{matrix}\right)+\left(\begin{matrix}n \\ 1\end{matrix}\right)+...+\left(\begin{matrix}n \\ n\end{matrix}\right)\] this one looks better

OpenStudy (anonymous):

can you use "pascal triangle" (the formula !)?

OpenStudy (anonymous):

wow never heard that~

OpenStudy (anonymous):

Your basis should be straight forward. I would start with your "n choose " side first. add n+1 by added a whole new term. You can then make a large chunk of it equal to 2^n. Now you just need to make that whole side look like 2^(n+1)

OpenStudy (experimentx):